OpenStudy (maemae16):

If p − 1 is a factor of p4 + p2 + p − k, the value of k is?

OpenStudy (maemae16):

@legomyego180

OpenStudy (phi):

if (p-1) is a factor , that means (p-1) divides evenly into the polynomial and we can write the polynomial as (p-1)* (other factors) if we set p=1 then (p-1)=0 and 0*(other factors) = 0 in other words, when p=1 we know p4 + p2 + p − k,= 0 so put in 1 for p, and then solve for k.